OpenStudy (haleholmes):

The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 times the height, what should each dimension be? Hint: Let x represent one of the dimensions, and then define the other dimensions in terms of x. if you could help me work it out that would be great. I really don't understand.

OpenStudy (phi):

calculus?

OpenStudy (haleholmes):

Algebra 2

OpenStudy (phi):

do you know the formula for the volume of a "box"?

OpenStudy (haleholmes):

L * w * h

OpenStudy (phi):

Let x represent one of the dimensions. which one do you want to call x ?

OpenStudy (haleholmes):

Um, height?

OpenStudy (phi):

or maybe, first analyze what we know. length must be 1.5 times the height can you write a formula that shows this ? Give it a try.

OpenStudy (haleholmes):

(1.5)L

OpenStudy (phi):

you should be writing an equation (which has an = sign in it) start by renaming the words with their "short names": length is L and height is H: length must be 1.5 times the height L is 1.5 * H can you change that into an equation ?

OpenStudy (haleholmes):

1.5H=L

OpenStudy (phi):

Yes, that is good.

OpenStudy (phi):

now let's tackle this part: sum of the dimensions at 6 inches can you write an equation for this?

OpenStudy (haleholmes):

L+W+H=6

OpenStudy (phi):

ok, good

OpenStudy (phi):

now if we call H "x" (yet a new name for height) we have H= x L= 1.5 H so what is L (if we use x instead of H ) ?

OpenStudy (haleholmes):

1.5x?

OpenStudy (phi):

yes, but get used to writing the whole equation (less chance of getting confused) L = 1.5 so here is what we have (so far): H= x L= 1.5 x H+L+W= 6 can you replace H and L in the last equation (H+L+W= 6) so we have x's ?

OpenStudy (phi):

**yes, but get used to writing the whole equation (less chance of getting confused) L = 1.5x

OpenStudy (haleholmes):

Okay, x+1.5x+W=6

OpenStudy (phi):

yes. can you solve for W ?

OpenStudy (haleholmes):

One sec

OpenStudy (haleholmes):